OpenStudy (anonymous):

Solve for x: 3|x - 3| + 2 = 14 No solutions x = -1, x = 8.3 x = 0, x = 7 x = -1, x = 7

OpenStudy (tester97):

\[3\left| x-3 \right|+2=14\] Subtract 2 frome both sides \[3\left| x-3 \right|=12\] Divide both sides by 3 \[\left| x-3 \right|=12\] can you do it from here?

OpenStudy (tester97):

Whoops i meant |x-3|=4

OpenStudy (anonymous):

No... Im lost completely... and im doing a test

OpenStudy (anonymous):

-7 and 7

OpenStudy (anonymous):

So D?

OpenStudy (tester97):

Yea it seems right but im not supposed to help you with tests......read the Code of conduct http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

So the answer is D..... Its like a pratice test

OpenStudy (tester97):

Go with what you believe is correct :D

OpenStudy (anonymous):

Please, im trying to get a good grade... Im begging... Is it D?

OpenStudy (tester97):

I gave you most of it now apply your knowledge you have and you will do fine ^_^

OpenStudy (anonymous):

Never mind -_-

OpenStudy (anonymous):

gosh please

OpenStudy (tester97):

I refuse to do your work for you so just apply yourself young man/woman i know you can!

OpenStudy (anonymous):

I know it can't be a zero, because that would result into no solution, so that leads to -1 and 7

OpenStudy (tester97):

Good job :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (tester97):

Welcome ^_~